I want to make sure that all of the work I have thus far is correct. And I need the answers to the analysisOn January 1, 2021, the general ledger of Big Blast Fireworks includes the following account balances: Accounts Cash AccountsCheck my work The following information is available on January 31, 2021. a. At the end of January, the company estimates thaNo Account Title Credit Date Jan 03 Debit 205,200 Inventory Accounts Payable 205,200 2 Jan 08 226,000 Inventory Accounts PayaCheck my work 8 Jan 24 Accounts Payable 650,000 Cash 650,000 Jan 27 3,500 Allowance for Uncollectible Accounts Accounts ReceiCheck my work January 31, 2021 Account Title Credit Cash $ Debit 134,900 68,000 5,240 15,000 76,600 38,700 Accounts ReceivablCheck my work Trial Balance January 31, 2021 Account Title Credit Debit 134,900 68,000 5,240 15,000 76,600 38,700 300 13,300LLLLL Choose the appropriate accounts to complete the companys income statement. The unadjusted, adjusted, or post-closing bAdjustea Big Blast Fireworks Classified Balance Sheet January 31, 2021 Liabilities Current Liabilities: $ 134,900 Accounts PaRequirement General Journal General Ledger Trial Balance Income Statement Balance Sheet Analysis Using the information from tCheck my work The inventory turnover ratio is: The company managing its inventory more efficiently. (True or False) (b) Calcu
Answer #1

a) Inventory turnover ratio = Cost of Goods Sold / Average Inventory
= $686300 / 27500 = 24.96
Average Inventory = ($40000+15000)/2 = $27500

The company is managing its inventory more efficiently, since ratio has increased

b) Gross Profit Ratio = Gross Profit / Sales x 100
= $228700/915000 = 24.99%

The company is less profitable

c) selling higher volume at lower prices